# Fedor V. Fomin

### v2

@@ 1.2 Overview of Theorems 1 and 2
Overview of the proof of Theorem 1: To sketch the proof of Theorem 1, we first recall the standard proof of the Gallai–Milgram theorem for undirected graphs. This proof hinges on a simple observation: if a path cover of $G$ consists of more than $\alpha(G)$ paths, then there is an edge connecting the endpoints of two different paths. Merging these two paths into one reduces the cover by one path. Reiterating this step yields a polynomial-time algorithm that produces a path cover of size at most $\alpha(G)$. Notably, the algorithm does not need to know $\alpha(G)$ in advance; it merely keeps merging pairs of paths with adjacent endpoints until no such pair remains. The procedure can begin from the trivial cover formed by $\lvert V(G)\rvert$ single-vertex paths.
The procedure above does not, by itself, provide any additional relationship between the number of paths $p$ in the obtained path cover $\mathcal{P}$ and the independence number $\alpha(G)$ beyond $p\leq\alpha(G)$. To address this issue, we introduce the notions of usual and special paths. A path is called usual if its endpoints are distinct and no edge connects them in $G$; otherwise, the path is special. We argue that having usual paths in $\mathcal{P}$ is advantageous. For instance, if all paths in $\mathcal{P}$ are usual, then the endpoints of these paths form an independent set of size $2p$. (Otherwise, we could merge two paths whose endpoints are adjacent.) In particular, this implies $p\leq\alpha(G)/2$. Even having $k$ usual paths is sufficient, since their endpoints would then form an independent set of size $p+k$. Alongside $\mathcal{P}$, this independent set would constitute a feasible output for the algorithm in Theorem 1.
@@
While special paths are not helpful for constructing a large enough independent set, they are convenient for merging. A special path is either a path of length zero or one, or it can be viewed (along with the edge between its endpoints) as a cycle of length at least three. Then, an edge between any two vertices of different special paths is as effective as an edge between the endpoints: it enables us to merge those two paths and cover all their vertices with a single path. Our algorithm takes advantage of this property by exhaustively merging special paths until no edges remain between them.
@@
This property of special paths offers additional insight into the structure of $G$.
Because there are no edges with endpoints in distinct special paths, and there are fewer than $k$ usual paths in $\mathcal{P}$, selecting one vertex from each special path already forms an independent set of size exceeding $p-k$. Moreover, if we choose two non-adjacent vertices from each of $2k$ special paths, we obtain an independent set of size at least $p+k$, immediately resolving the problem. Thus, the only remaining scenario is that, except for at most $2k$ special paths, vertex set of each of the special paths is a clique in $G$.
However, this structural insight alone addresses only part of the problem. To progress further, we require a more sophisticated transformation of the paths in $\mathcal{P}$. Rather than merging two paths, we now consider two special paths and one usual path such that there are two edges (with no shared endpoints) connecting the special paths to the usual path. Using these edges, the transformation creates three new paths, at least two of which become usual. As a result, the number of special paths in $\mathcal{P}$ decreases without increasing the total number of paths (see Figure 1).
If all these transformations cannot be applied, we arrive at the following situation. There must exist a vertex set $S$ of size smaller than $k$ that separates all usual paths from at least $p-4k$ special paths, each of which induces a clique in $G$. In the auxiliary Lemma 1, we demonstrate how this small separator $S$ can be exploited to remove irrelevant cliques from $G$ without affecting the size of an optimal path cover or the independence number. Applying this polynomial-time procedure ensures that, in the end, at most $\mathcal{O}(k^{2})$ special paths remain in $\mathcal{P}$.
By applying transformations and irrelevant cliques removals, we end up in the situation, where $p=\mathcal{O}(k^{2})$. In this case, we invoke Theorem 2 to compute an optimal path cover of $G$, using $p+k$ as its parameter. If Theorem 2 does not return an optimal path cover, it instead provides an independent set of size $p+k$. In either outcome, the algorithm obtains a valid solution for Theorem 1.
@@
Overview of the proof of Theorem 2: We focus on outlining the main ideas of the proof specifically for $-\ell$-linkages. The more general extension to arbitrary list TM-embeddings is similar but requires additional technical details.
@@
To begin, we sketch an $\operatorname{{\sf X}P}$ algorithm that solves $-\ell$-Linkage in time $n^{f(\alpha(G)+\ell)}$. Although this algorithm does not achieve the desired complexity, it is instructive for understanding the proof. Suppose we are given a parameter $k\geq\alpha(G)$. (This assumption is acceptable for an $\operatorname{{\sf X}P}$ algorithm because $k$ can be computed by brute force in $\mathcal{O}(n^{\alpha(G)})$ time.) The $\operatorname{{\sf X}P}$ algorithm is recursive and is inspired by the Chvátal-Erdős’s theorem on highly connected graphs, see Proposition 1. Informally, Chvátal-Erdős’s theorem states that if a graph lacks a large independent set but is highly connected, it necessarily contains a Hamiltonian path. If its connectivity is low, we find a small vertex cut, delete it, and then recursively process the connected components of the resulting graph. Because a Hamiltonian path might traverse each component multiple times, we find it more convenient to work with $-\ell$-linkages, that is, linkages consisting of vertex-disjoint paths connecting specified pairs of vertices and covering all vertices of the graph. To handle such cases, one can strengthen Chvátal-Erdős theorem to apply to Hamiltonian-linkages. Following the foundational work of Bollobás and Thomason [BT96], there have been numerous advances in understanding how linkages relate to connectivity. However, our setting additionally requires that the linkages are Hamiltonian. While no existing results directly address this requirement, we adapt the result of Thomas and Wollan [TW05] on linkages in highly connected graphs to show that if a graph $G$ is $\Omega(\alpha(G)\cdot\ell)$-vertex-connected, then it is $-\ell$-linked. In other words, there exists a path cover of size $\ell$ that links any prescribed set of endpoints.
@@
The algorithm proceeds as follows. If the input graph $G$ is $\Omega(k\cdot\ell)$-connected, then the algorithm correctly reports that $G$ is $-\ell$-linked. Otherwise, if the connectivity is lower, there exists a minimum vertex separator $X$ of $G$ with $\mathcal{O}(k\cdot\ell)$ vertices. Each connected component of $G-X$ has an independence number at most $\alpha(G)-1$, and the algorithm proceeds recursively on each component. For that, the algorithm performs an exhaustive search of how the solution paths are split into parts by $X$. For every connected component of $G-X$ and each split, the algorithm also performs an exhaustive search of entry and exit points of the parts of the paths in the component. Since the number of parts is at most $\ell+2|X|$, this constitutes $n^{\mathcal{O}(\ell+|X|)}$ possible choices. The algorithm then treats these points as new terminals and makes a recursive call for each connected component. Restricting the recursion depth by $\alpha(G)$, the recursion reaches the subproblem where the graph has sufficiently high connectivity without significantly blowing up the number of paths. Overall, this yields a running time of $n^{f(\alpha(G)+\ell)}$.
@@
Since the $\operatorname{{\sf X}P}$ algorithm exploits small separators separating highly connected components, the first natural idea that comes to mind is to check whether the generic tools about cuts from the toolbox of parameterized algorithms could be leveraged to turn it into an $\operatorname{{\sf FPT}}$ algorithm. The recursive understanding technique introduced by Chitnis et al. [CCH+16] is arguably the most powerful tool for designing $\operatorname{{\sf FPT}}$ algorithms for such types of problems about cuts, as illustrated by the meta-theorem of Lokshtanov et al. [LRSZ18]. Unfortunately, recursive understanding does not appear to be particularly helpful in our setting for the following reason: in order to apply recursive understanding, one must handle the so-called $(p,q)$-unbreakable graphs as the “base case.” However, for our problem, addressing $(p,q)$-unbreakable graphs is not simpler than handling the general case.
@@
Thus, we need a different method of decomposing the graph. To prevent the $\operatorname{{\sf X}P}$ running time, the $\operatorname{{\sf FPT}}$ algorithm avoids recursive calls to itself. Instead, it encapsulates the recursion in a subroutine that in time $f(k,\ell)\cdot n^{\mathcal{O}(1)}$ either output an independent set of $G$ of size $k$ or find a vertex set $S$ of size $k^{\mathcal{O}(k)}\cdot\ell$ with the following property: Each connected component of $G-S$ is $\Omega(k\cdot(\ell+|S|))$-connected. Set $S$ breaks solution paths into at most $\ell+2|S|$ parts. Some parts belong to $G[S]$, while others belong to $G-S$. Since the size of $S$ is in $k^{\mathcal{O}(k)}\cdot\ell$, the paths in $G[S]$ could be enumerated in $\operatorname{{\sf FPT}}$ (in $k$ and $\ell$) time. However, guessing the entry and exit points of these parts, the way we describe it for the $\operatorname{{\sf X}P}$ algorithm, is too time-consuming. Here is precisely the moment when the property of the decomposition comes into play. Each connected component of $G-S$ is highly connected and thus is Hamiltonian-$(\ell+2|S|)$-linked. Therefore, for any selection of the paths’ endpoints, the required collection of vertex-disjoint paths spanning all component vertices always exists (unless empty). Therefore, it is sufficient for the algorithm to enumerate all possible “patterns” of paths’ behavior between $S$ and the components of $G-S$ without enumerating the entry-exit endpoints in $G-S$. The latter task reduces to computing a maximum matching in an auxiliary graph. This way, the whole enumeration can be done in $\operatorname{{\sf FPT}}$ time.
@@
Another complication arises because we do not assume $\alpha(G)<k$. (Recall that we cannot compute $\alpha(G)$ in $\operatorname{{\sf FPT}}$ time.) Instead, we aim either to find a desired solution or an independent set of size $k$. This requirement calls for a constructive variant of the high-connectivity case. Specifically, given an $\Omega(k\cdot\ell)$-connected graph and $\ell$ pairs of endpoints, we must be able to construct $\ell$ disjoint paths connecting these endpoints and covering all vertices of the graph, or else find an independent set of size $k$. For this purpose, we rely on a combinatorial result by Thomas and Wollan [TW05], which states that in a $10\ell$-connected graph, one can always find $\ell$ vertex-disjoint paths connecting any given pairs of endpoints. However, the proof in [TW05] is non-constructive—it builds on certain properties of a minimal counterexample, and it is not obvious how to transform that argument into an algorithm that yields the required paths. To circumvent this issue, we make use of the fact that our graphs have bounded independence number. Through a Ramsey-based argument, we either locate an independent set of size $k$ or find a large clique. In the latter scenario, we can apply Menger’s theorem to connect the endpoints to the clique, thereby constructing the required $\ell$ paths.
@@
As a side remark, several combinatorial results—including the aforementioned work of Thomas and Wollan [TW05]—guarantee the existence of $\ell$ vertex-disjoint paths between specified pairs of endpoints in a graph with connectivity $\Omega(\ell)$. However, there is a notable shortage of simple and efficient algorithms to compute such vertex-disjoint paths in $\Omega(\ell)$-connected (or even $\ell^{\Omega(1)}$-connected) graphs. Our contribution in this direction may therefore be of independent interest.
@@
Finally, for Theorem 2, we additionally generalize the argument above from $-\ell$-linkages to arbitrary list TM-embeddings. Here, once more, the non-recursive structure of the algorithm proves helpful: With the same choice of the set $S$, we enumerate all possible behaviors of the target model on $S$ and between $S$ and the connected components of $G-S$. Again, we avoid enumerating concrete terminals and connecting vertices in $G-S$, fixing only the general connection pattern to $S$. While considerably more technical, the enumeration can still be done in $\operatorname{{\sf FPT}}$ time. Then, similarly to $-\ell$-linkages, it can be shown that one can always find a spanning list TM-model in a highly-connecting component (or an independent set of size $k$), with arbitrary locations of the terminals.
1.3 Related Work.

### v3

@@ 1.2 Overview of Theorems 1 and 2
Overview of the proof of Theorem 1: To sketch the proof of Theorem 1, we first recall the standard argument for Gallai–Milgram’s theorem in undirected graphs. The key observation is that if a path cover of $G$ has more than $\alpha(G)$ paths, then there must be an edge connecting the endpoints of two distinct paths. By merging these paths, the cover size decreases by one. Thus, starting from the trivial cover consisting of $\lvert V(G)\rvert$ single-vertex paths, we can iteratively obtain a path cover of size at most $\alpha(G)$. This procedure runs in polynomial time. Importantly, we do not need to know $\alpha(G)$ in advance; the algorithm simply keeps merging paths with adjacent endpoints until no such pair remains.
The merging procedure used to prove Gallai–Milgram’s theorem does not provide any additional information beyond $p\leq\alpha(G)$, where $p$ is the size of the constructed path cover $\mathcal{P}$. To move beyond this bound, we extract more structural information from a path cover once the Gallai–Milgram merging process can no longer continue. For this purpose, we introduce the notions of usual and special paths. A path is called usual if its endpoints are distinct and no edge connects them in $G$; otherwise, the path is special.
Intuitively, usual paths are “open-ended”: their endpoints are free, and since no edge ties them together, both endpoints can safely be included in an independent set. Usual paths are therefore useful for building independence, as each such path potentially contributes two vertices to the independent set. More formally, assume there are at least $k$ usual paths in $\mathcal{P}$ and that no edges connect the endpoints of two distinct paths $P_{i},P_{j}\in\mathcal{P}$. Then, if for each special path (of which there are at most $|\mathcal{P}|-k=p-k$) we take one endpoint, and for each usual path we take both endpoints, we obtain an independent set $I$ of size at least $(p-k)+2k\geq p+k$. The path cover $\mathcal{P}$ together with the independent set $I$ forms a certificate in Theorem 1 for the case when $\mathcal{P}$ contains at most $\alpha(G)-k$ paths.
Thus, the first step of the proof of Theorem 1 is as follows. We start from a trivial path cover $\mathcal{P}$ consisting of single-vertex paths and perform “Gallai–Milgram mergings”. That is, we merge paths in $\mathcal{P}$ exhaustively as long as there exists a pair of paths with adjacent endpoints. When no such merge is possible, if at least $k$ paths are usual, the algorithm constructs $I$ and outputs a valid solution.
In the remainder of the proof, the algorithm considers the case when in the constructed path cover $\mathcal{P}$:
  • (i)
    No two endpoints of distinct paths are adjacent, and
  • (ii)
    fewer than $k$ paths are usual, so that the remaining paths are special.
@@
Informally, special paths are “closed”: they either consist of a single vertex (a trivial path), or have two vertices, or are of length at least two and their endpoints are connected by an edge, so that the path together with this edge forms a cycle. In this situation, an edge between any two vertices of different special paths plays the same role as an edge between their endpoints: it allows us to merge the two paths and cover all their vertices with a single path. Therefore, while special paths are not very useful for constructing independent sets, they are extremely convenient for merging. Our algorithm exploits this property by exhaustively merging special paths until no edges remain between them.
As a result, we obtain the following structural property of the path cover:
  • (iii)
    if two distinct paths $P_{i},P_{j}\in\mathcal{P}$ are special, then there is no edge between $V(P_{i})$ and $V(P_{j})$ in $G$.
We perform a similar merging trick in the case when $P_{i}$ is special, $P_{j}$ is usual, and there is an edge from $V(P_{i})$ to an endpoint of $P_{j}$. This yields the fourth property:
  • (iv)
    If a path $P_{i}\in\mathcal{P}$ is special, then there is no edge between $V(P_{i})$ and any endpoint of any other path $P_{j}\in\mathcal{P}$.
Together, properties (i)–(iv) provide additional structural insight into the path cover $\mathcal{P}$ and, consequently, into the graph $G$. Armed up with these properties, our algorithm is now able to construct a larger independent set in $G$ than before. For each usual path, the algorithm still takes its both endpoints into $I$. For each special path $P_{i}\in\mathcal{P}$, the algorithm can now take two arbitrary non-adjacent vertices from $V(P_{i})$ into $I$ provided that such vertices exist. If $V(P_{i})$ is instead a clique in $G$, the algorithm takes just a single arbitrary vertex into $I$. The resulting set $I$ is independent by properties (i), (iii), (iv) of $\mathcal{P}$. If at least $k$ special paths do not induce a clique in $G$, the size of $I$ is at least $p+k$, and this immediately resolves the problem. Thus, the only remaining scenario is that,
  • (v)
    except for at most $k$ special paths, the vertex set of each of the special paths is a clique in $G$.
@@
However, this structural insight alone addresses only a part of the problem. To progress further, we need a more sophisticated transformation of the paths in $\mathcal{P}$. Rather than merging two paths, we now consider two special paths and one usual path such that there are two edges (with no shared endpoints) connecting the special paths to the usual path. Using these edges, the transformation creates three new paths, at least two of which become usual. As a result, the number of special paths in $\mathcal{P}$ decreases without increasing the total number of paths (see Figure 1).
If all these transformations cannot be applied, we arrive at the following situation. There must exist a vertex set $S$ of size smaller than $k$ that separates all usual paths from at least p4kp-4k special paths, each of which induces a clique in $G$. In the auxiliary Lemma 1, we demonstrate how this small separator $S$ can be exploited to remove irrelevant cliques from $G$ without affecting the size of an optimal path cover or the independence number. Applying this polynomial-time procedure ensures that, in the end, at most $\mathcal{O}(k^{2})$ special paths remain in $\mathcal{P}$.
By applying transformations and irrelevant clique removals, we end up in the situation where $p=\mathcal{O}(k^{2})$. In this case, we invoke Theorem 2 to compute an optimal path cover of $G$, using $p+k$ as its parameter. If Theorem 2 does not return an optimal path cover, it instead provides an independent set of size $p+k$. In either outcome, the algorithm obtains a valid solution for Theorem 1.
@@
Overview of the proof of Theorem 2: We focus on outlining the main ideas of the proof specifically for Hamiltonian-$\ell$-linkages. The more general extension to arbitrary list TM-embeddings is similar but requires additional technical details.
@@
To begin, we sketch an $\operatorname{{\sf X}P}$ algorithm that solves Hamiltonian-$\ell$-Linkage in time $n^{f(\alpha(G)+\ell)}$. Although this algorithm does not achieve the desired complexity, it is instructive for understanding the proof. Suppose we are given a parameter $k\geq\alpha(G)$. (This assumption is acceptable for an $\operatorname{{\sf X}P}$ algorithm because $k$ can be computed by brute force in $\mathcal{O}(n^{\alpha(G)})$ time.) The $\operatorname{{\sf X}P}$ algorithm is recursive and is inspired by the Chvátal-Erdős’s theorem on highly connected graphs, see Proposition 1. Informally, Chvátal-Erdős’s theorem states that if a graph lacks a large independent set but is highly connected, it necessarily contains a Hamiltonian path. If its connectivity is low, we find a small vertex cut, delete it, and then recursively process the connected components of the resulting graph. Because a Hamiltonian path might traverse each component multiple times, we find it more convenient to work with Hamiltonian-$\ell$-linkages, that is, linkages consisting of vertex-disjoint paths connecting specified pairs of vertices and covering all vertices of the graph. To handle such cases, one can strengthen Chvátal-Erdős’s theorem to apply to Hamiltonian-linkages. Following the foundational work of Bollobás and Thomason [14], there have been numerous advances in understanding how linkages relate to connectivity. However, our setting additionally requires that the linkages are Hamiltonian. While no existing results directly address this requirement, we adapt the result of Thomas and Wollan [87] on linkages in highly connected graphs to show that if a graph $G$ is $\Omega(\alpha(G)\cdot\ell)$-vertex-connected, then it is Hamiltonian-$\ell$-linked. In other words, there exists a path cover of size $\ell$ that links any prescribed set of endpoints.
@@
The algorithm proceeds as follows. If the input graph $G$ is $\Omega(k\cdot\ell)$-connected, then the algorithm correctly reports that $G$ is Hamiltonian-$\ell$-linked. Otherwise, if the connectivity is lower, there exists a minimum vertex separator $X$ of $G$ with $\mathcal{O}(k\cdot\ell)$ vertices. Each connected component of GXG-X has an independence number at most $\alpha(G)-1$, and the algorithm proceeds recursively on each component. For that, the algorithm performs an exhaustive search of how the solution paths are split into parts by $X$. For every connected component of GXG-X and each split, the algorithm also performs an exhaustive search of entry and exit points of the parts of the paths in the component. Since the number of parts is at most $\ell+2|X|$, this constitutes $n^{\mathcal{O}(\ell+|X|)}$ possible choices. The algorithm then treats these points as new terminals and makes a recursive call for each connected component. Restricting the recursion depth by $\alpha(G)$, the recursion reaches the subproblem where the graph has sufficiently high connectivity without significantly blowing up the number of paths. Overall, this yields a running time of $n^{f(\alpha(G)+\ell)}$.
@@
Since the $\operatorname{{\sf X}P}$ algorithm exploits small separators separating highly connected components, the first natural idea that comes to mind is to check whether the generic tools about cuts from the toolbox of parameterized algorithms could be leveraged to turn it into an $\operatorname{{\sf FPT}}$ algorithm. The recursive understanding technique introduced by Chitnis et al. [17] is arguably the most powerful tool for designing $\operatorname{{\sf FPT}}$ algorithms for such types of problems about cuts, as illustrated by the meta-theorem of Lokshtanov et al. [75]. Unfortunately, recursive understanding does not appear to be particularly helpful in our setting for the following reason: in order to apply recursive understanding, one must handle the so-called $(p,q)$-unbreakable graphs as the “base case.” However, for our problem, addressing $(p,q)$-unbreakable graphs is not simpler than handling the general case.
@@
Thus, we need a different method of decomposing the graph. To prevent the $\operatorname{{\sf X}P}$ running time, the $\operatorname{{\sf FPT}}$ algorithm avoids recursive calls to itself. Instead, it encapsulates the recursion in a subroutine that in time $f(k,\ell)\cdot n^{\mathcal{O}(1)}$ either outputs an independent set of $G$ of size $k$ or finds a vertex set $S$ of size $k^{\mathcal{O}(k)}\cdot\ell$ with the following property: Each connected component of GSG-S is $\Omega(k\cdot(\ell+|S|))$-connected. Set $S$ breaks solution paths into at most $\ell+2|S|$ parts. Some parts belong to $G[S]$, while others belong to GSG-S. Since the size of $S$ is in $k^{\mathcal{O}(k)}\cdot\ell$, the paths in $G[S]$ could be enumerated in $\operatorname{{\sf FPT}}$ (in $k$ and $\ell$) time. However, guessing the entry and exit points of these parts, the way we describe it for the $\operatorname{{\sf X}P}$ algorithm, is too time-consuming. Here is precisely the moment when the property of the decomposition comes into play. Each connected component of GSG-S is highly connected and thus is Hamiltonian-$(\ell+2|S|)$-linked. Therefore, for any selection of the paths’ endpoints, the required collection of vertex-disjoint paths spanning all component vertices always exists (unless empty). Therefore, it is sufficient for the algorithm to enumerate all possible “patterns” of paths’ behavior between $S$ and the components of GSG-S without enumerating the entry-exit endpoints in GSG-S. The latter task reduces to computing a maximum matching in an auxiliary graph. This way, the whole enumeration can be done in $\operatorname{{\sf FPT}}$ time.
@@
Another complication arises because we do not assume $\alpha(G)<k$. (Recall that we cannot compute $\alpha(G)$ in $\operatorname{{\sf FPT}}$ time.) Instead, we aim either to find a desired solution or an independent set of size $k$. This requirement calls for a constructive variant of the high-connectivity case. Specifically, given an $\Omega(k\cdot\ell)$-connected graph and $\ell$ pairs of endpoints, we must be able to construct $\ell$ disjoint paths connecting these endpoints and covering all vertices of the graph, or else find an independent set of size $k$. For this purpose, we rely on a combinatorial result by Thomas and Wollan [87], which states that in a $10\ell$-connected graph, one can always find $\ell$ vertex-disjoint paths connecting any given pairs of endpoints. However, the proof in [87] is non-constructive—it builds on certain properties of a minimal counterexample, and it is not obvious how to transform that argument into an algorithm that yields the required paths. To circumvent this issue, we make use of the fact that our graphs have bounded independence number. Through a Ramsey-based argument, we either locate an independent set of size $k$ or find a large clique. In the latter scenario, we can apply Menger’s theorem to connect the endpoints to the clique, thereby constructing the required $\ell$ paths.
@@
As a side remark, several combinatorial results—including the aforementioned work of Thomas and Wollan [87]—guarantee the existence of $\ell$ vertex-disjoint paths between specified pairs of endpoints in a graph with connectivity $\Omega(\ell)$. However, there is a notable shortage of simple and efficient algorithms to compute such vertex-disjoint paths in $\Omega(\ell)$-connected (or even $\ell^{\Omega(1)}$-connected) graphs. Our contribution in this direction may therefore be of independent interest.
@@
Finally, for Theorem 2, we additionally generalize the argument above from Hamiltonian-$\ell$-linkages to arbitrary list TM-embeddings. Here, once more, the non-recursive structure of the algorithm proves helpful: With the same choice of the set $S$, we enumerate all possible behaviors of the target model on $S$ and between $S$ and the connected components of GSG-S. Again, we avoid enumerating concrete terminals and connecting vertices in GSG-S, fixing only the general connection pattern to $S$. While considerably more technical, the enumeration can still be done in $\operatorname{{\sf FPT}}$ time. Then, similarly to Hamiltonian-$\ell$-linkages, it can be shown that one can always find a spanning list TM-model in a highly connected component (or an independent set of size $k$), with arbitrary locations of the terminals.
1.3 Related Work